A piece of string whose length is 32cm is cut into 2 pieces. one piece is used to form an equilateral triangle and the other to form of the circle so that the sum of the areas is a minimum? Find the minimum sum of the areas. Express your answer in pi. What is the function









Computation of derivative









Computation of Critical numbers









Computation of minimum value










1
Expert's answer
2022-03-21T06:43:36-0400

Solution.

Determine:

x cm - a piece is used to form an equilateral triangle.

32-x cm - a piece is used to form a circle.

"S_1=\frac{a^2\sqrt 3}{4}" - the area of the triangle.

"a=\frac x3" ; "S_1=\frac{x^2\sqrt 3}{36}".

"S_2=\pi r^2" - the area of the circle.

"l=2\pi r" ; "r=\frac {l}{2\pi}=\frac{32-x}{2\pi}";

"S_2=\pi (\frac{32-x}{2\pi})^2=\frac {1}{4\pi}(32-x)^2".

The sum of the areas:

"S=S_1+S_2= \frac{x^2\sqrt 3}{36}+ \frac {1}{4\pi}(32-x)^2" .

Derivative of the function:

"S’=\frac{2x\sqrt 3}{36}-\frac{32-x}{2\pi}=0"

"\frac{2x\sqrt 3}{36}=\frac{32-x}{2\pi}"

"\frac{ x\sqrt 3}{9}=\frac{32-x}{\pi}"

"x_0=\frac{9\cdot32}{\pi\sqrt 3+9}" - critical number.

As "S’’=\frac {\sqrt 3}{18}+\frac {1}{2\pi}>0" so "S(x_0)" is the minimum value of the function "S(x)".

"S(x_0)=(\frac{9\cdot 32}{\pi\sqrt 3+9})^2\cdot\frac{\sqrt 3}{36}+\frac{1}{4\pi}(32-(\frac{9\cdot 32}{\pi\sqrt 3+9})^2)^2="

"\frac{256}{\pi+3\sqrt 3}"

Answer: "S_{min}= \frac{256}{\pi+3\sqrt 3}" "cm^2" .